OpenStudy (anonymous):

6/17+2/8 lowest terms

jimthompson5910 (jim_thompson5910):

in order to add these two fractions, the denominators must be the same

jimthompson5910 (jim_thompson5910):

the LCM of 17 and 8 is 17*8 = 136 so the common denominator is 136

jimthompson5910 (jim_thompson5910):

multiply top and bottom of the first fraction by 8 6/17 = (6*8)/(17*8) = 48/136 so... 6/17 = 48/136

jimthompson5910 (jim_thompson5910):

multiply top and bottom of the second fraction by 17 2/8 = (2*17)/(8*17) = 34/136 so... 2/8 = 34/136

jimthompson5910 (jim_thompson5910):

so put this all together to get 6/17+2/8 to be the same as 48/136 + 34/136 can you take it from here?